Determine the volume of  C2H5OH(l) , in mL , that was combusted to produce the volume of  CO2(g)  collected in part (b). (The density of  C2H5OH(l)  is 0.79g/mL .)
1
Expert's answer
2020-05-03T15:10:35-0400

Molar mass of Co2 is 44.01g/mol and mass of co2 is 15

"Number of moles = mass\/molar mass."

Moles of Co2=15/44.01= 0.3408mol

From the equation ,1 mol of C2H5OH =0.5 mol of Co2. Therefore, 0.1704moles of C2H5OH are present.

Molar mass of C2H5OH is 46.07g/mol and its mass is 0.1704*46.07=7.851 . volume will be given by mass/density=

9.95ml